OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 14 −20 36 −3 −11

OpenStudy (bahrom7893):

14 = -2x + 8 7 = -x +4 7-4 = -x 3=-x x = -3